OpenStudy (anonymous):

Solve the equation by extracting square roots. 3x^2 = 81

OpenStudy (anonymous):

3x^2 = 81 x^2 = 27 x = root 27, - root 27 x = 3 root 3, - 3 root 3

OpenStudy (anonymous):

\[3x^2=81\] \[x^2=27\] \[x=\pm\sqrt{27}=\pm3\sqrt{3}\]